Martin Manufacturing has implemented several programs to improve its productivity. They have asked you to evaluate the firm's productivity by comparing this year's performance with last year's. The following data are available:

Last Year This Year
Output 10,500 units 12,100 units
Labor Hours 12,000 13,200
Utilities $7,600 $8,250
Capital $83,000 $88,000

a. What is the percent change in labor hours this year vs. last year?                            [ Select ]                       ["7.69%", "3.25%", "5.34%", "4.76%"]      

b. What is the percent change in utility cost this year vs. last year?                            [ Select ]                       ["6.52%", "-1.25%", "4.55%", "6.16%"]      

c. What is the percent change in capital investment this year vs. last?                            [ Select ]                       ["9.24%", "-6.75%", "5.55%", "8.69%"]      

d. What is the overall multi-factor productivity for last year?                            [ Select ]                       [".01267", ".10233", ".2457", ".11055"]      

e. What is the overall multi-factor productivity this year?                            [ Select ]                       [".04677", ".34780", ".11055", ".10233"]      

f. What is the multi-factor percent change this year vs. last year?                            [ Select ]                       ["-2.7%", "8.0%", "3.5%", "11.8%"]      
Answer #1

PRODUCTIVITY = TOTAL OUTPUT / TOTAL INPUT

CHANGE IN PRODUCTIVITY = NEW PRODUCTIVITY - OLD PRODUCTIVITY / OLD PRODUCTIVITY * 100

PARTIAL PRODUCTIVITY

PRODUCTIVITY

LAST

CURRENT

% CHANGE

LABOR

0.875

0.9167

4.77

UTILITIES

1.3816

1.4667

6.16

CAPITAL

0.1265

0.1375

8.7

MULTIFACTOR PRODUCTIVITY

LAST YEAR

VALUES:

LABOR = 12000

UTILITIES = 7600

CAPITAL = 83000


TOTAL OUTPUT = 10500

PRODUCTIVITY = TOTAL OUTPUT / TOTAL INPUT

PRODUCTIVITY = 10500 / (12000 + 7600 + 83000) = 0.10234


THIS YEAR

VALUES:

LABOR = 13200

UTILITIES = 8250

CAPITAL = 88000


TOTAL OUTPUT = 12100

PRODUCTIVITY = TOTAL OUTPUT / TOTAL INPUT

PRODUCTIVITY = 12100 / (13200 + 8250 + 88000) = 0.11055


% CHANGE

CHANGE IN PRODUCTIVITY = NEW PRODUCTIVITY - OLD PRODUCTIVITY / OLD PRODUCTIVITY * 100


PRODUCTIVITY = ((0.11055 - 0.10234) / 0.10234) * 100 = 8.02228 OR 8.0

TABLE


LAST YEAR

0.10234

THIS YEAR

0.11055

CHANGE

8


**PLEASE LEAVE A THUMBS UP